OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

jimthompson5910 (jim_thompson5910):

what is f(3)?

OpenStudy (anonymous):

i think it 9 but I'm not sure

jimthompson5910 (jim_thompson5910):

let's find out. f(3) happens when n = 3 f(n) = f(1) + f(2) + f(n - 1) f(3) = f(1) + f(2) + f(3 - 1) ... replace all copies of n with 3 f(3) = f(1) + f(2) + f(2) f(3) = 2 + 5 + 5 f(3) = 12

jimthompson5910 (jim_thompson5910):

see how I got 12?

OpenStudy (anonymous):

yea so... whats next, thats where i get kind of lost

jimthompson5910 (jim_thompson5910):

what is f(4)

OpenStudy (anonymous):

10?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

f(4) happens when n = 4 f(n) = f(1) + f(2) + f(n - 1) f(4) = f(1) + f(2) + f(4 - 1) ... replace all copies of n with 4 f(4) = f(1) + f(2) + f(3) f(4) = ??

OpenStudy (anonymous):

oh so its 22?

jimthompson5910 (jim_thompson5910):

what is f(1)

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

what is f(2)

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

what is f(3)

OpenStudy (anonymous):

12

jimthompson5910 (jim_thompson5910):

f(4) happens when n = 4 f(n) = f(1) + f(2) + f(n - 1) f(4) = f(1) + f(2) + f(4 - 1) ... replace all copies of n with 4 f(4) = f(1) + f(2) + f(3) f(4) = 2 + 5 + 12 f(4) = 19

OpenStudy (anonymous):

ok so the f(5)=F(!)+f(2)+f(3)+f(4)

jimthompson5910 (jim_thompson5910):

no you follow the definition f(n) = f(1) + f(2) + f(n - 1)

jimthompson5910 (jim_thompson5910):

and plug in n = 5

OpenStudy (anonymous):

so f(5)=11

jimthompson5910 (jim_thompson5910):

no f(5) has to be bigger than f(4)

jimthompson5910 (jim_thompson5910):

f(5) happens when n = 5 f(n) = f(1) + f(2) + f(n - 1) f(5) = f(1) + f(2) + f(5 - 1) ... replace all copies of n with 5

OpenStudy (anonymous):

ok so it would look like this f(5) = 2+ 5 + (5 - 1)

jimthompson5910 (jim_thompson5910):

f(5) = 2 + 5 + f(5 - 1)

jimthompson5910 (jim_thompson5910):

i see how you got 11 now

jimthompson5910 (jim_thompson5910):

that 5-1 turns into 4, but you cannot add it to 7 to get 11

jimthompson5910 (jim_thompson5910):

f(5) happens when n = 5 f(n) = f(1) + f(2) + f(n - 1) f(5) = f(1) + f(2) + f(5 - 1) ... replace all copies of n with 5 f(5) = f(1) + f(2) + f(4)

jimthompson5910 (jim_thompson5910):

you need to replace f(4) with what we just found f(4) to be

OpenStudy (anonymous):

f(5)=2+5+19 f(5)=26

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thx

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

help me with another

jimthompson5910 (jim_thompson5910):

sure, i can do one more

OpenStudy (anonymous):

Use the graph below to fill in the blank with the correct number. f(0) = _______

jimthompson5910 (jim_thompson5910):

f(0) means "f(x) when x = 0"

jimthompson5910 (jim_thompson5910):

so you're looking for a point with an x coordinate of zero

jimthompson5910 (jim_thompson5910):

f(0) will be the point (0, y) where y = f(0)

OpenStudy (anonymous):

so ..

jimthompson5910 (jim_thompson5910):

any ideas where this point is?

OpenStudy (anonymous):

(0,3)

OpenStudy (anonymous):

sorry (3,0))

jimthompson5910 (jim_thompson5910):

(0,3) or (3,0) is not a point plotted

OpenStudy (anonymous):

(-3,0))

jimthompson5910 (jim_thompson5910):

how many points are on this graph?